how fast must a cyclist climb a 6.0 degrees hill to maintain a power output of 0.2 hp? Neglect work done by friction and assume the mass of cyclists and bicycle is 68 kg
1
Expert's answer
2020-11-23T10:28:09-0500
"P=Fv=mgv\sin{6}\\0.2(745.7)=(98)(9.8)\sin{6}v\\v=1.5\frac{m}{s}"